### v2

# Discovery of the Hidden World with Large Language Models

Abstract
Revealing the underlying causal mechanisms in the real world is the key to the development of science. Despite the progress in the past decades, traditional causal discovery approaches (CDs) mainly rely on high-quality measured variables, usually given by human experts, to find causal relations. The lack of well-defined high-level variables in many real-world applications has already been a longstanding roadblock to a broader application of CDs. To this end, this paper presents Causal representatiOn AssistanT (COAT) that introduces large language models (LLMs) to bridge the gap. LLMs are trained on massive observations of the world and have demonstrated great capability in extracting key information from unstructured data. Therefore, it is natural to employ LLMs to assist with proposing useful high-level factors and crafting their measurements. Meanwhile, COAT also adopts CDs to find causal relations among the identified variables as well as to provide feedback to LLMs to iteratively refine the proposed factors. We show that LLMs and CDs are mutually beneficial and the constructed feedback provably also helps with the factor proposal. We construct and curate several synthetic and real-world benchmarks including analysis of human reviews and diagnosis of neuropathic and brain tumors, to comprehensively evaluate COAT. Extensive empirical results confirm the effectiveness and reliability of COAT with significant improvements.
1 Introduction
Science originates along with identifying important variables and revealing their causal relations [1, 2]. Despite the progress in the past decades, existing causal discovery approaches (CDs) mainly rely on high-quality measured variables, which are usually given by human experts [3, 4, 5]. However, the desired variables and their measurements are usually unavailable in a wide range of real-world applications. For example, Amazon sellers who want to analyze the factors related to user ratings only have raw user reviews, which are written according to the underlying user preferences for certain product characteristics. Therefore, the lack of high-quality high-level variables has been a longstanding impediment to broader real-world applications of CDs or causality-inspired methods [6].
Figure 1: Illustration of COAT framework to analyze the rating scores of AppleGastronome. COAT aims to uncover the underlying Markov Blanket with respect to the given ratings of the apples (i.e., factors that fit the preferences of gastronomes). COAT first (a) adopts an LLM to read, comprehend, and relate the rich knowledge about tasting the apples. The LLM needs to propose a series of candidate factors such as apple sizes and smells, along with some meta-information such as annotation guidelines. Based on the candidate factors, COAT then (b) prompts another LLM to annotate the unstructured review into structured data. (c) The CD algorithm then finds causal relations among the factors, and constructs feedback based on samples where the ratings can not be well explained by the existing factors. By looking into the new samples, the LLM is expected to associate more related knowledge to uncover more desired causal factors.
The recently emerged Large Language Models (LLMs) [7, 8, 9, 10] offer a new opportunity to mitigate the gap [11]. Trained from massive observations of the world, LLMs demonstrate impressive capabilities in comprehending unstructured inputs, and leveraging the learned rich knowledge to resolve a variety of general tasks [12]. A surge of early tests demonstrates promising results that LLMs can effectively leverage the learned knowledge to answer commonsense causal questions [13, 11, 14]. Nevertheless, existing approaches mainly focus on incorporating LLMs as a straightforward reasoner with respect to the given causal variables. The reliability of LLMs in directly reasoning the true causal structure behind any specific data-generating process remains a debate [15, 16, 17, 13] due to a series of drawbacks of LLMs [18, 19, 20]. In addition, all of the existing combinations of LLMs and causal discovery have surprisingly overlooked the identifiability of causal structure, which plays an important role in classic causal discovery literature [3, 4, 5]. Hence, a challenging question comes:
How can LLMs reliably assist in revealing the causal mechanisms behind the real world?
In this work, we answer the question with a focus on local causal discovery with respect to a target variable that poses high value such as customer ratings and medical diagnosis, and introduce Causal representatiOn AssistanT (COAT). Specifically, given the target variable $Y$, COAT aims to identify a Markov blanket to $Y$ from raw observation and also produce the theoretical-guaranteed causal structure. To achieve the goal, as illustrated in Fig. 1, COAT employs two mutually beneficial components: LLMs and CDs. Iteratively, at step (a), COAT leverages LLMs to look into a set of unstructured observations (e.g., customer comments) and propose potentially useful high-level factors. The proposed factors contain both the definition and the annotation criteria. Therefore, at step (b), another LLM is employed to give concrete values following the criteria. Then in step (c), CDs is used to reveal the structure among the identified factors. To ensure the reliability of the factor identification, COAT constructs feedback from the intermediate causal discovery results from step (c) to further inspire LLMs to improve further factors. The feedback includes sampling important observations that can not be well explained by the existing identified results. We show that the feedback provably helps with identifying the desired Markov Blanket and the structure (Proposition 2.2). We present a comprehensive analysis of COAT on both synthetic simulations and real-world case studies, ranging from analysis of human reviews to diagnosis of neuropathic and brain tumors (Sec. 4). Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to leverage LLMs to propose high-level variables, thereby extending the scope of CDs to unstructured data (Sec. 2.1).
  • We establish the first benchmarks with real-world data including AppleGastronome and Neuropathic to examine the unstructured causal discovery (Sec. 4).
  • We propose the first framework COAT that combines the best of LLMs and CDs to find theoretically grounded causal results (Sec. 2.2), which are verified with extensive empirical studies (Sec. 4).
  • Additionally, the analysis of COAT also derives the first metrics that measure the causal representation learning capabilities of various LLMs (Definition 2.3).
2 Representation Assistant for Causal Discovery
In this section, we present the formulation of leveraging a language model to serve as a representation assistant for causal discovery on unstructured data. The representation assistant needs to extract useful factors that capture sufficient information for an interested target variable.
2.1 Problem Definition
Data    To begin with, we are given a target variable of interest $Y$, e.g., stars rated by a customer or the tumor type of a patient. We treat $Y$ as a scalar random variable without loss of generality. The unstructured data, or raw observations, e.g., customer review of a certain product or images of tumors, are denoted as ${\bm{X}}$. The dataset ${\mathcal{D}}$ consists of $n$ paired samples $\left\{\left({\bm{x}}_{i},y_{i}\right)\right\}_{i=1}^{n}$ that are independently drawn from the distribution over $({\bm{X}},Y)$. Note that the target variable $Y$ serves as a guider, and no prior assumption on the relations between ${\bm{X}}$ and $Y$ is assumed.
Objective    We seek a mapping $h:{\bm{X}}\mapsto{\bm{Z}}$ which elicits the structured representation ${\bm{Z}}=h({\bm{X}})$ such that $Y\perp\!\!\!\!\perp{\bm{X}}\mid{\bm{Z}}$. In other words, ${\bm{Z}}$ serves as a Markov Blanket of $Y$ for the unstructured raw observations. Built upon the structure representation, then, downstream causal discovery methods can be applied on ${\bm{Z}}\cup\{Y\}$. The revealed causal structure can provide insights about the target variable $Y$ [21, 22], such as what factors of the product would be most satisfactory to customers. Furthermore, the framework can be easily extended to discover a complete causal graph by shifting the target variable to the other identified factors or the other additionally available variables. Formal assumptions are discussed in Sec. 2.4.
LLM as a representation assistant    We aim to make the most use of the rich knowledge of LLMs to assist in extracting the relevant information from the raw observations ${\bm{X}}$. To this end, the mapping $h$ is decomposed as a collection of factors ${\mathcal{W}}=\left\{{\bm{w}}_{1},{\bm{w}}_{2},...,{\bm{w}}_{k}\right\}$, each of which is a function ${\bm{w}}_{i}:{\bm{X}}\mapsto{\mathcal{C}}$ that maps the raw observation ${\bm{x}}$ to a predefined value space ${\mathcal{C}}$. In other words, the structured representation is composed of multiple factors: $h({\bm{X}})=\left({\bm{w}}_{1}\left({\bm{X}}\right),{\bm{w}}_{2}\left({\bm{X}}\right),\cdots,{\bm{w}}_{k}\left({\bm{X}}\right)\right)$. Throughout this work, for the notation of factors, we use the symbol ${\bm{w}}_{i}$ to denote the factor itself like sweetness, size, or scent, and ${\bm{w}}_{i}(\cdot)$ to denote the function that maps from raw observation space ${\mathcal{X}}$ to the predefined value space ${\mathcal{C}}$.
Descriptions of the factors    Without loss of generality, in this work, we consider the descriptions of the factors in natural language, which can be divided into two categories: i) Implicit factors, which need to be discovered and elaborated by LLMs. To obtain the values of the implicit factors, one could feed the factor descriptions and the unstructured input ${\bm{x}}_{i}$ to a suitable LLM for value annotation; For example, given a customer review on an apple ${\bm{x}}_{i}$, a discovered implicit factor ${\mathcal{C}}=\left\{-1,0,1\right\}$: ${\bm{w}}_{1}({\bm{x}}_{i})=1$ could mean that the customer appreciates the sweetness of the apple; ${\bm{w}}_{1}({\bm{x}})=-1$ means that the customer is disappointed about the sweetness; ${\bm{w}}_{1}({\bm{x}})=0$ means the sweetness has not been mentioned. ii) Explicit factors such as heart rate, whose descriptions are already available. The measure of the explicit factors usually requires some external tools.
2.2 The COAT Framework
We approach the aforementioned problem via a new framework called Causal representatiOn AssistanT (COAT) (Algorithm 1). COAT aims to extract useful factors through multiple rounds of iteration. We use the superscript $t$ to refer to the input and output of LLM at the $t$-th round. We also denote the union of results from the first $t$ rounds using the superscript “$\leq t$”.
Algorithm 1 The COAT Framework
1:  Required: Dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$; LLM for factor proposal $\Psi$; Model for factor parsing $\Psi_{s}$; causal discovery algorithm ${\mathcal{A}}$; Feedback constructor ${\mathcal{F}}$; Maximal rounds $T$;
2:  Random sampling $\widehat{{\mathcal{D}}}^{1}$;
3:  Constructing ${\bm{p}}^{1}$;
4:  while not converge and current round $t<T$ do
5:     ${\mathcal{W}}^{t}\leftarrow\Psi({\bm{p}}^{t},\widehat{{\mathcal{D}}}^{t})$; //factor proposal
6:     ${{\bm{Z}}}^{t}\leftarrow\Psi_{s}({\mathcal{D}},{{\mathcal{W}}}^{t},{\bm{p}}_{p})$; //factor parsing
7:     ${{\mathcal{G}}}^{t}\leftarrow{\mathcal{A}}({{\bm{Z}}}^{\leq t}\cup\{Y\})$; //causal discovery
8:     $(\widehat{{\mathcal{D}}}^{t+1},{\bm{p}}^{t+1})\leftarrow{\mathcal{F}}({{\mathcal{G}}}^{t},{\mathcal{D}},{\bm{p}}^{t})$; //feedback
9:  end while
10:  return ${\mathcal{G}}^{T}$
Factor proposal
To induce useful high-level factors from the rich world knowledge of LLMs, COAT employs a prompt ${\bm{p}}$ and a few samples $\widehat{{\mathcal{D}}}\subseteq{\mathcal{D}}$.
Specifically, in Fig. 2, ${\bm{p}}$ contains three components: samples, instructions, and format control. To encourage LLM to focus on the information related to the target variable $Y$, samples are grouped by the value of $Y$. The instruction requires $\Psi$ to give each proposed factor ${\bm{w}}_{i}$ a concrete description of the mapping ${\bm{w}}_{i}(\cdot)$, like how to decide the factor values. In addition, the metadata about the task such as the task description and context can also be incorporated if available. The prompt ${\bm{p}}$ essentially imitates human experts [23] in selecting and defining high-level variables. The set of resulting factors in the $t$-th round, defined with natural language by the LLM $\Psi$, is denoted as ${\mathcal{W}}^{t}=\Psi({\bm{p}}^{t},\widehat{{\mathcal{D}}}^{t})$. We merge them with factors proposed in the previous rounds to update the set of all factors ${\mathcal{W}}^{\leq t}={\mathcal{W}}^{1}\cup\cdots\cup{\mathcal{W}}^{t}$.
Figure 2: Illustration of the prompt template for factor proposal in COAT.
Factor parsing
Once we obtain the candidate factors, we then collect the values of the factors from the unstructured observations. In prior works, they are usually collected from human experts according to the given factors [3]. To do so, another LLM $\Psi_{s}$ is instructed to read the annotated guidelines of each variable ${\bm{w}}_{i}$ and parse the unstructured observations into structured or tabular data:
${{\bm{z}}}_{i}:=\big{(}{\bm{w}}_{i}({\bm{x}}_{1}),\cdots,{\bm{w}}_{i}({\bm{x}}_{n})\big{)},\quad{\bm{w}}_{i}({\bm{x}}):=\Psi_{s}({\bm{x}},{\bm{w}}_{i},{\bm{p}}_{p}),$ (1)
where ${\bm{p}}_{p}$ refers to the additional instruction to prompt $\Psi_{s}$ to parse the observed data, and ${{\bm{z}}}_{i}$ refers to the parsed values for the corresponding factor $\widehat{{\bm{w}}}_{i}$. We define ${{\bm{Z}}}^{\leq t}:=\text{Concat}\big{(}\{{{\bm{z}}}_{i}\,\text{for}\,{\bm{w}}_{i}\in{\mathcal{W}}^{\leq t}\}\big{)}$.
When the data curation of the proposed factors requires additional domain-specific knowledge/skills (e.g., intervening on the external environments) that the LLMs do not acquire, we could fetch ${{\bm{z}}}_{i}$ through some external process [24, 25]. For example, studying the causes of a disease requires annotating relevant symptoms from diagnosis records and conducting additional medical checks  [26]. Our experiments show that COAT can effectively extract the hidden factors under both schemes.
Causal discovery
With the given values ${{\bm{Z}}}^{\leq t}$ associated with the candidate factors ${\mathcal{W}}^{\leq t}$, a CD algorithm ${\mathcal{A}}$ (e.g., FCI [27]) is used to reason about the causal structure based on the parsed data:
${{\mathcal{G}}}^{t}={\mathcal{A}}({{\bm{Z}}}^{\leq t}\cup\{Y\}),$ (2)
where ${{\mathcal{G}}}^{t}$ is the discovered causal structure. In general, the inputs in each round to ${\mathcal{A}}$ may contain noises as well as latent confounders, any CDs with suitable theoretical assumptions could be used for ${\mathcal{A}}$. The noises injected through LLM-based parsing may be of independent interest to the literature of causal discovery [28, 5]. In this work, we demonstrate the idea of COAT via the FCI algorithm [4] as it is flexible with respect to different functional classes of the underlying generation process, allows for the existence of latent confounders [27], which aligns well with our objective.
Improving factor proposal with causal feedback
LLMs require proper prompts to fully unlock their capabilities [29, 30, 31]. When it comes to factor proposing, it is also hard for LLMs to propose all factors at once. Nevertheless, from the causal discovery results, we could find useful information and thus provide feedback to further improve the factor proposal:
$(\widehat{{\mathcal{D}}}^{t+1},{\bm{p}}^{t+1})={\mathcal{F}}({{\mathcal{G}}}^{t},{\mathcal{D}},{\bm{p}}^{t}),$ (3)
where ${\mathcal{F}}$ samples specific examples from ${\mathcal{D}}$ and constructs new prompts according to the results of ${\mathcal{A}}$ for the next round of factor proposal. For example, FCI is able to imply the existence of latent confounders, from which we could refine ${\bm{p}}$ to prompt $\Psi$ to focus on the corresponding factors.
Figure 3: Illustration of variables that could be discovered with COAT. Let $W\in h_{\leq t}(X)$ be an identified variable, and assume there exists a latent variable $\widehat{{\bm{w}}}$ to be discovered. When $\widehat{{\bm{w}}}$ is the direct parent or child of $Y$, finding hard-to-explain samples can help uncover it. When $\widehat{{\bm{w}}}$ is the direct parent and also a child of $W$, or the spouse of $Y$ with $W$ as the common child of $Y$ and $\widehat{{\bm{w}}}$, conditioning on $W$ facilitates the discovery of $\widehat{{\bm{w}}}$.
2.3 Causal Feedback
[…]
$H(Y|h_{\leq t}(X))>H(Y|h_{\leq t}(X),\widehat{{\bm{w}}}(X)),$ (4)
where $H(\cdot)$ refers to the entropy. If the LLM $\Psi$ can not find the desired $\widehat{{\bm{w}}}$, it means that $h_{\leq t}(X)$ is already sufficient to separate $Y$ from $X$. Therefore, for the next $(t+1)$-th iteration with sample $\widehat{{\mathcal{D}}}^{t+1}$, $\Psi$ is expected to propose new factor $\widehat{{\bm{w}}}$ that also satisfies the similar property:
$H_{\widehat{{\mathcal{D}}}^{t+1}}(Y|h_{\leq t}(X))-H_{\widehat{{\mathcal{D}}}^{t+1}}(Y|h_{\leq t}(X),\widehat{{\bm{w}}}(X))>0,$ (5)
where $H_{\widehat{{\mathcal{D}}}^{t}}(\cdot|\cdot)$ refers to the conditional entropy measured on $\widehat{{\mathcal{D}}}^{t}$. As shown in Fig. 3, finding factors satisfying Eq. 4 progressively expands the discovered factors and pushes $h_{\leq t}(X)$ to a valid Markov Blanket. Also, it implies conditioning on the identified factors would further strengthen the correlation between $\widehat{w}$ and $Y$. Therefore, to find the desired factor, we are motivated to select suitable $\widehat{{\mathcal{D}}}^{t+1}$ for the next iteration such that $\widehat{{\mathcal{D}}}^{t+1}=\text{$\operatorname*{arg\,max}$}_{\widehat{{\mathcal{D}}}\subseteq{\mathcal{D}}}H_{\widehat{{\mathcal{D}}}}(Y|h_{\leq t}(X)),$ where $h_{\leq t}(X)$ can not well explain $Y$. This problem can be converted into a classification problem in which $\widehat{{\mathcal{D}}}^{*}$ are the samples that the fitted classifier yields a large prediction error. In our experiments, we implement the classification via clustering with respect to $h_{\leq t}(X)$. The clustering elicits $C$ groups $\widehat{{\mathcal{D}}}_{c}:=\big{\{}{\bm{x}}_{i}\,\,\text{for}\,\,i\in{\mathcal{I}}_{c}\big{\}}$: ${\mathcal{I}}_{1},\cdots,{\mathcal{I}}_{C}=\text{K-Means}\big{(}h_{\leq t}(X)\big{)}$. We then take the group of samples with the largest conditional entropy to construct the feedback.
In practice, many factors, such as the LLM capabilities, data faithfulness, and prompt templates, could affect the satisfaction of Eq 5. Therefore, in the next section, we will establish a theoretical framework to discuss the influence of the factors above to the satisfaction of Eq 5.
2.4 Theoretical Analysis
We then theoretically analyze some critical steps in COAT, including the feedback and identifiability.
Feedback analysis    Given a new factor ${\bm{w}}_{k+1}$, with the current representation as $h_{[k]}(X)=\left({\bm{w}}_{1}\left(X\right),{\bm{w}}_{2}\left(X\right),\cdots,{\bm{w}}_{k}\left(X\right)\right)$, and COAT tests:
[…]
COAT also requires the following usual condition about distribution and causal graph:
Assumption 2.1 (Faithful and Markov conditions, adjusted from  [33]).
For any disjoint non-empty subsets $A,B,C\subset\mathcal{W}^{\leq t}\cup\{Y\}$, $A$ and $B$ are d-separated by $C$ on the causal graph iif $A\perp\!\!\!\!\perp B\mid C$ on the factors’ distribution. All conditional independencies are preserved after factor parsing.
The annotation from a poor model could introduce an additional “error term” on the true factor values, disturbing the true distribution, as one can observe in Fig. 4(a) and 4(b). If Assumption 2.1 holds, the conditional mutual information between $Y$ and $X$ given the desired factors decreases:
Proposition 2.2.
Under assumption 2.1, if condition 6 holds, then for Markov Blanket ${\mathcal{S}}\subseteq[k+1]$ of $Y$, i.e., $Y\perp\!\!\!\!\perp h_{[k+1]\setminus{\mathcal{S}}}(X)\mid h_{{\mathcal{S}}}(X)$, we have the following about conditional mutual information:
$I\left(Y;X\mid h_{{\mathcal{S}}}(X)\right)=I\left(Y;X\mid h_{[k+1]}(X)\right)<I\left(Y;X\mid h_{[k]}(X)\right)$ (7)
Factor identification    We provide an initial exploration under what conditions LLMs can identify target-related factors and how the ability of an LLM influences this procedure.
Definition 2.3 (Ability of LLMs).
Given a suitable prompt about current factors and data, the LLM $\Psi$ has non-zero probability $p_{\Psi}>0$ to propose a new factor ${\bm{w}}_{k+1}$ that satisfies condition 6 and
$\frac{I\left(Y;X\mid h_{[k+1]}(X)\right)}{I\left(Y;X\mid h_{[k]}(X)\right)}<1-C_{\Psi},$ (8)
for some positive constant $C_{\Psi}$ whenever $I\left(Y;X\mid h_{[k]}(X)\right)>0$. Note that $h_{[0]}(X)=\phi$, hence we also use $I\left(Y;X\mid h_{[0]}(X)\right)$ to refer $I\left(Y;X\right)$. We use $p$ instead of $p_{\Psi}$ when the context is clear.
We further explain the intuition behind Def 2.3: the Perception Score $p$ captures the LLM’s responsiveness to the given prompts and the feedback; the Capacity Score $C_{\Psi}$ captures the quality of the factors proposed by the LLM. Empirically, the two scores are used to estimate the abilities of the predominant LLMs (Sec. 3.2). Theoretically, we use them to characterize the influence of prompt templates, the LLM responsiveness, and the quality of factors on the performance of COAT:
Proposition 2.4 (Characterization for Factor Identification Process).
With assumption 2.1, for any small number $\epsilon,\delta\in(0,\frac{1}{2})$, perception score $p>0$, capacity score $C_{\Psi}>0$, with $t$ COAT rounds that
$\sqrt{t}>\frac{|z_{\delta}|\sqrt{1-p}}{2\sqrt{p}}\left(1+\sqrt{1+\frac{4\log{\epsilon}}{z_{\delta}^{2}(1-p)\log{(1-C_{\Psi})}}}\,\right),$ (9)
where $z_{\delta}$ is the $\delta$-quantile of the standard normal distribution, we have
$\Pr\left(\frac{I\left(Y;X\mid h_{\leq t}(X)\right)}{I\left(Y;X\right)}<\epsilon\right)\geq 1-\delta.$ (10)
The proof is given in Appendix D.2. Prop. 2.4 gives a guarantee on identifying a Markov Blanket. Intuitively, Prop. 2.4 also characterizes the influence of prompt templates, the LLM responsiveness, and the quality of factors on the performance of COAT via the two proposed measures: $p$ and $C_{\Psi}$. When both of them are positive, COAT would converge exponentially:
Proposition 2.5 (Rate of Convergence).
With assumption 2.1, for any small number $\epsilon,\delta\in(0,\frac{1}{2})$, perception score $p>0$, capacity score $C_{\Psi}>0$, after $t$ COAT rounds, the following inequality holds with probability at least $1-\delta$:
$\frac{I(Y;X\mid h_{\leq t}(X))}{I(Y;X)}\leq\left(\frac{1}{1-C_{\Psi}}\right)^{-tp-z_{\delta}\sqrt{tp(1-p)}}$ (11)
Causal structure identification    It is clear that LLMs are not involved in the causal discovery process, which is mainly executed by causal discovery methods such as FCI. Therefore, the CD guarantees the identifiability of the final causal graph over the LLM-proposed factors. The concrete assumptions required for identifiability depend on the specific CD used in COAT. For instance, the FCI algorithm requires faithfulness of the data distribution with respect to the true causal graph [27]. In our experiments, we also verify that the structured data annotated by LLMs has a high accuracy and little noise, which is friendly to the CD assumption. In general, one could switch to another CD in COAT, while using different CDs may require different assumptions. For example, the LiNGAM algorithm requires the relations among variables to be linear and non-Gaussian models. Empirically, we find that COAT with LiNGAM works very well (Appendix F).
2.5 Practical Discussions
Prompt template    The instruction following capability and the context window of LLMs may affect the satisfaction to the constraints of the prompt template. Including more data samples or background knowledge may improve the $p$ and $C_{\Psi}$, but it is more challenging for the LLM.
Modern causal discovery    We use FCI in this paper in order to illustrate the idea. To attain identifiability better than the Markov equivalent class, one can choose more advanced Causal Discovery methods under different assumptions, see Appendix C if interested.
We also discuss some cases where we need to handle them properly in practice with LLMs.
Factor filtering    LLMs may output several factors with similar semantics or exhibit multicollinearity in the annotated data, which will hinder the causal discovery process. To mitigate the issue, one could do factor filtering that adopts PCA or early conditional independence tests given the currently discovered variables in the Markov Blanket to detect and eliminate these variables.
Factor pool    LLMs may discover useful factors in early rounds while being discarded. For example, the underlying spouse variables of the target label $y$ may be independent with $y$ without conditioning on their common children variables. To resolve the issue, we could introduce a factor pool that stores the candidate variables proposed in the past, and replay the variables that have not been passed by conditional independence tests with existing variables in the Markov Blanket for a double check.
3 Empirical Analysis of COAT
We evaluate whether COAT can propose and identify a set of high-level factors belonging to the Markov Blanket of the target variable $Y$. We construct the first benchmark, called AppleGastronome, to verify the effectiveness of COAT in finding useful causal information, and compare COAT to previous methods [11] that merely leverage LLMs to perform causal discovery. Specifically, we use AppleGastronome to examine the capabilities of $10$ predominant LLMs including GPT 4o [34], Claude-3-Opus [35], LLaMA3-70b [36], and Mistral-Large [37] in realizing COAT. Due to the space limit, we report only the results of the popular LLMs and present full results in Appendix E.4.
(a) Apple attributes Acc.
(b) Preference matchness Acc.
(c) LLM Ability Scatters
Figure 4: Quantitative evaluation of the causal capabilities of LLMs in COAT.
3.1 Experimental Setup
In AppleGastronome benchmark, we consider the target variable as a rating score of the apple by astronomers. We prepare different high-level factors: 3 parents of Y, one child of Y, and one spouse of Y. These factors form a Markov blanket of Y. In addition, we also prepared one disturbing factor related to Y but not a part of this blanket. A good method is expected to propose the five high-level factors (up to semantical meanings) and exclude the disturbing factor.
Benchmark construction    Each apple has its own attributes, including size, smell, and taste (or sweetness). Each gastronomy pays unique attention to a subset of the above three attributes. They will write a review according to their preference and give the rating score. We generate the review using GPT 4 by feeding GPT 4 the preferences and the apple attributes. We generated 200 samples for LLMs’ analysis and annotation. More details are left in Appendix E.1.
Baselines    For factor proposal, we mainly employ two different uses of LLMs as the baselines: META is the zero-shot factor proposal given only the context to LLMs; and DATA additionally gives some samples of raw observations, which is an ablation of COAT without the feedback module, i.e., only one COAT round. For causal relation inference, we follow Kiciman et al. [11] that prompt LLMs to reason for the causal direction of each pair of the discovered variables by DATA.
Metrics We evaluate the ability on factor proposal based on three metrics: MB, NMB, and OT. MB means the desired factor forming the Markov Blanket of Y. NMB means the undesired factor relevant to data but not in MB. OT means the unexpected factors irrelevant to data. We also present the corresponding recall, precision, and F1 with respect to $\text{MB}(Y)$.
(a) Ground truth
(b) GPT-4 META
(c) GPT 3.5 COAT
(d) Claude-3-Opus COAT
Figure 5: The discovered causal graphs in AppleGastronome. Compared to the ground truth results, directly adopting LLMs to reason the causal relations can easily elicit many false positive edges. In contrast, the relations recovered by COAT have a high precision and recall. The directed edge between “taste” and “juiciness” can not be recovered by COAT because of the limitations of FCI.
3.2 Analysis with AppleGastronome Benchmark
Table 1: Factor proposal results in Apple Gastronome benchmark (Full results in Appendix E.4).
LLM Method Factor Proposal
MB NMB OT Recall Precision F1
GPT 4 META 2.67$\pm$0.94 0.67$\pm$0.47 2.33$\pm$0.47 0.53$\pm$0.19 0.46$\pm$0.08 0.49$\pm$0.13
DATA 3.00$\pm$0.00 0.33$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.92$\pm$0.12 0.72$\pm$0.04
DATA+CoT 4.33$\pm$0.58 0.83$\pm$0.29 0.17$\pm$0.29 0.87$\pm$0.12 0.81$\pm$0.02 0.84$\pm$0.06
COAT 4.00$\pm$0.82 0.33$\pm$0.47 0.00$\pm$0.00 0.80$\pm$0.16 0.93$\pm$0.09 0.85$\pm$0.11
GPT 3.5 META 3.33$\pm$1.25 0.33$\pm$0.47 4.33$\pm$1.25 0.67$\pm$0.25 0.42$\pm$0.12 0.51$\pm$0.15
DATA 2.67$\pm$0.47 0.67$\pm$0.47 0.00$\pm$0.00 0.53$\pm$0.09 0.81$\pm$0.14 0.64$\pm$0.10
DATA+CoT 5.00$\pm$0.00 1.00$\pm$0.00 1.33$\pm$0.58 1.00$\pm$0.00 0.68$\pm$0.05 0.81$\pm$0.04
COAT 3.67$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.73$\pm$0.09 1.00$\pm$0.00 0.84$\pm$0.07
LLaMA2 -70b META 2.33$\pm$0.47 0.67$\pm$0.47 4.67$\pm$1.25 0.47$\pm$0.09 0.32$\pm$0.07 0.37$\pm$0.06
DATA 2.33$\pm$0.94 0.67$\pm$0.47 0.00$\pm$0.00 0.47$\pm$0.19 0.75$\pm$0.20 0.57$\pm$0.20
DATA+CoT 3.00$\pm$1.73 0.67$\pm$0.58 0.33$\pm$0.58 0.60$\pm$0.35 0.71$\pm$0.34 0.65$\pm$0.35
COAT 3.00$\pm$0.00 0.67$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.83$\pm$0.12 0.69$\pm$0.04
Mistral -MEDIUM META 3.00$\pm$0.00 0.67$\pm$0.47 1.67$\pm$1.25 0.60$\pm$0.00 0.59$\pm$0.13 0.59$\pm$0.07
DATA 3.00$\pm$0.00 0.67$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.83$\pm$0.12 0.69$\pm$0.04
DATA+CoT 4.33$\pm$0.58 1.00$\pm$0.00 0.67$\pm$0.58 0.87$\pm$0.12 0.73$\pm$0.07 0.79$\pm$0.05
COAT 4.67$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.93$\pm$0.09 1.00$\pm$0.00 0.96$\pm$0.05
Key findings    Empirically, LLMs with CoT can be aware of high-level factors behind data (lower OT than META) but still struggle to distinguish the desired factors in Markov Blanket (higher NMB than COAT). COAT is more resistant to the “disturbing” factor, which is supported by the lower NMB column. COAT filters out irrelevant factors from LLMs’ prior knowledge that are not reflected by the data, which is supported by the lower OT column. COAT robustly encourages LLM to find more expected factors through the feedback, which is supported by the higher MB column.
Can LLMs be an effective factor proposer?    As discussed in Sec. 2.4, there are two crucial abilities for LLMs in identifying potential high-level factors. The first one is to be aware of the existence of potential factors, and the second is to synthesize and describe these factors. Inspired by this observation, we propose two novel metrics to quantify LLMs’ causal ability: a) Perception that quantifies the ratio of valid factors (satisfying Prop. 2.2) proposed by LLMs in each round; b) Capacity that measures the effective mutual information drop in Assumption 2.3. As shown in Fig. 4(c), LLMs differ largely on the perception score while comparably on the capacity score.
Can LLMs be an effective factor annotator?    As shown in Fig. 4, both GPT 3.5 and GPT 4 annotate subjective attributes well. Regarding objective human preferences, the performances are still relatively high. Empirically, LLMs will not introduce new confounders, see Appendix E.3.
Can COAT reliably recover the causal relationships?    We present quantitative and qualitative results in Table 3 (in Appendix E.1) and Fig. 5 (on page 8), respectively. Compared to directly adopting LLMs to reason the causal relations, COAT significantly boosts the causal relation recovery. Meanwhile, COAT maintains high performances based on various LLMs, which further demonstrates the effectiveness of the causal feedback in COAT to improve the robustness of this system. In fact, the causal feedback focuses on making maximal use of the rich knowledge of LLMs, and reducing the reliance on the reasoning capabilities of different LLMs, to assist with causal discovery. We provide the full results of $10$ LLMs in Appendix E.3.
(a) Faithful ground truth
(b) GPT-4 COAT
(c) GPT-4 reasoning
Figure 6: The discovered causal graphs in Neuropathic. (c) shows the result based on directly prompting LLM to reason for the causal relations among all factors. Disconnected ones are dropped.
Table 2: Factor proposal results in Neuropathic. PA, AN, and OT refer to the parents, ancestors, and others, respectively. Accuracy and F1 measure the recovery of the causal ancestors.
LLM Method Factor Proposal
PA AN OT Acc F1
GPT 4 Meta 3 5 6 0.91 0.59
Data 2 2 0 0.95 0.50
DATA+CoT 3 4 13 0.81 0.35
COAT 3 6 3 0.96 0.80
GPT 3.5 Meta 3 5 6 0.91 0.59
Data 3 5 4 0.94 0.67
DATA+CoT 2 2 3 0.91 0.36
COAT 3 5 2 0.96 0.77
LLaMA2 -70b Meta 2 4 5 0.91 0.53
Data 3 3 1 0.95 0.60
DATA+CoT 2 4 7 0.88 0.47
COAT 3 6 2 0.97 0.86
LLaMA2 -13b Meta 1 3 6 0.88 0.40
Data 3 6 4 0.95 0.75
DATA+CoT 0 1 10 0.81 0.12
COAT 3 6 2 0.97 0.86
LLaMA2 -7b Meta 1 1 17 0.72 0.08
Data 3 6 3 0.96 0.80
DATA+CoT 0 0 10 0.79 --
COAT 3 6 2 0.97 0.86
Mistral -Medium Meta 3 6 3 0.96 0.80
Data 3 3 2 0.94 0.66
DATA+CoT 3 5 8 0.88 0.53
COAT 3 6 2 0.97 0.86
4 Empirical Study with Realistic Benchmarks
After examining the capabilities of COAT in AppleGastronome, we are further motivated to challenge COAT in a more complex setting from neuropathic panic diagnosis [26], brain tumor detection with MRI images [38], three-years news summary about one stock from the New York Times [39], and climatic reanalysis data with fine-grained time and space coverage [40]. We refer to Appendix H for a complete summary of all five benchmarks.
4.1 Experimental Setup
Benchmark construction    In the Neuropathic benchmark, we convert the dataset into a clinical diagnosis task. In the original dataset, there are three levels of causal variables, including the symptom level, radiculopathy level, and the pathophysiology level. In the experiments, we mainly consider the target variable of right shoulder impingement. When generating the clinical diagnosis notes as ${\bm{x}}$ using GPT 4, we avoid any mentioning of variables other than symptoms. We generated 100 samples for LLMs’ analysis; since the number of possible factors is finite, we generate 1000 tabular data for CI tests.
As we intend to leverage the Neuropathic benchmark to simulate the real-world diagnosis, after the factor proposal stage, we directly incorporate external tools to measure the values of the candidate factors. More details about the construction of the Neuropathic are given in Appendix E.6.
Evaluation and baselines    In Neuropathic, we adopt a similar evaluation protocol and the baselines as in AppleGastronome. Nevertheless, due to the faithfulness issue of the original dataset [26], for the evaluation of causal relation discovery, we mainly conduct a qualitative comparison between the ground truth that is faithful to the data, against the baselines and COAT.
4.2 Empirical Results on Neuropathic Benchmark
Factor proposal    The quantitative results on Neuropathic benchmark are given in Table LABEL:table:neuropathic. Similarly, we can find that COAT consistently outperforms all of the baselines regardless of which LLMs are incorporated. In particular, even with the weakest backbone model, i.e.,LLaMA2-7b, COAT can still effectively leverage the intrinsic rich knowledge and beat the baselines with more powerful LLMs.
Causal relation recovery    Fig. 6(a) shows the causal graph obtained by FCI running on the original data, where we can find that several causal relations cannot hold on the data. As shown in Fig. 6, when using LLMs to perform the reasoning, LLMs cannot identify the faithfulness issues. In contrast, COAT can imply faithful causal insights.
Figure 7: The final causal graph found by COAT in the ENSO case study
4.3 More Real-world Results
El Niño–Southern Oscillation (ENSO) case study    El Niño-Southern Oscillation (ENSO) is a climatic phenomenon in the Pacific Ocean that influences global weather patterns profoundly. To understand its mechanism, we apply COAT on NOAA dataset [40]. There are 13 factors identified by COAT, and their instantaneous causal relations are visualized in Fig 7. The target variable is the future change in monthly SST in the Nino3 region, which could be an important indicator of ENSO events. Each factor is a time series about a certain climate measurement above a specific level averaged over a specific region. The paths about Sea level Pressure, Momentum Flux, and Cloud Cover matches the existing understanding from literature [41, 42, 43, 44]. It also suggests several inserting hypotheses that are less explored in literature, like the path from Soil Temperature in South American Coastal Region. We refer details in Appendix K.
More real-world empirical studies    We also report concrete results on real-world problems involving MRI, time series, and NetCDF data in Appendix IJ, and K.
5 Conclusions
In this paper, we proposed a new paradigm COAT to incorporate the rich knowledge of LLMs into the CD pipeline. We found that COAT effectively extends the scope of CDs to unstructured data by identifying useful high-level variables from raw observations for CD methods. COAT suggests a new pathway towards building a causal foundation model for discovery. We leave more detailed discussions about future studies in Appendix B.
Acknowledgments and Disclosure of Funding
We thank the reviewers for their valuable comments. This material is based upon work supported by NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Salesforce, Apple Inc., Quris AI, and Florin Court Capital. CXL and BH were supported by NSFC General Program No. 62376235, Guangdong Basic and Applied Basic Research Foundation Nos. 2022A1515011652 and 2024A1515012399, HKBU Faculty Niche Research Areas No. RC-FNRA-IG/22-23/SCI/04, and HKBU CSD Departmental Incentive Scheme. MMG was partially supported by the following Australian Research Council projects: DE210101624 and DP240102088. JC was supported by CUHK direct grant 4055146.
Appendix of COAT
[…]
Appendix A Table of Notations
Notation Description
$\Psi$ the LLM for factor proposal.
${\bm{p}}^{t}$ Prompt used at the $t$-th round for factor proposal.
$\Psi_{p}$ the LLM for factor parsing.
${\bm{p}}_{p}$ Prompt used for factor parsing.
${\mathcal{A}}$ Algorithm for causal discovery. In this paper we mainly use FCI.
${\mathcal{F}}$ Abstraction of the COAT’s Feedback procedure.
${\mathcal{D}}$ Full sample set.
$\widehat{{\mathcal{D}}}$ A subset of sample set
$\widehat{{\mathcal{D}}}^{t}$ The subset of sample used at the $t$-th round.
${\mathcal{C}}$ The predefined value space for LLM-proposed factors, e.g., set ${\mathcal{C}}$ as $\{-1,0,1\}$
${\bm{x}}$ An unstructured sample of raw observation drawn from ${\mathcal{X}}$
$y$ A structured sample of target variable drawn from ${\mathcal{Y}}$
$Y$ target variable treated as a random variable.
$X$ raw observation treated as a random variable.
${\bm{w}}_{i}$ The $i$-th factor proposed by LLMs, like sweetness, size, or scent
${\bm{w}}_{i}(\cdot)$ Factor’s corresponding mapping from raw observation space ${\mathcal{X}}$ to the predefined value space ${\mathcal{C}}$.
${\mathcal{W}}^{t}$ The set of factors proposed by LLMs at the $t$-th round.
${\mathcal{W}}^{\leq t}$ The set of factors proposed by LLMs at the first $t$ rounds.
${\mathcal{S}}$ Index set of selected factors in ${\mathcal{W}}$, e.g., ${\mathcal{S}}=\{1,3\}$ means only select the first and third factors.
$h_{{\mathcal{S}}}(\cdot)$ LLM-induced representation over selection ${\mathcal{S}}$, e.g., $h_{\{1,3\}}(\cdot)=\left({\bm{w}}_{1}(\cdot),{\bm{w}}_{3}(\cdot)\right)$. $h(\cdot):=h_{[k]}(\cdot)$
${{\bm{Z}}}$ Matrix of values of the structured high-level factors proposed by LLMs, i.e., $\widehat{{\bm{Z}}}\in{\mathcal{C}}^{n\times k}$
${{\bm{Z}}}^{t}$ Matrix of values of the structured high-level factors proposed at the $t$-th COAT round.
${{\bm{Z}}}^{\leq t}$ Matrix of values of the structured high-level factors proposed at the first $t$ COAT rounds.
${{\mathcal{G}}}^{t}$ Causal graph at the $t$-th COAT round over $\widehat{{\bm{Z}}}^{\leq t}\cup\{Y\}$
$k$ Number of factors
$n$ Number of samples
$C$ Number of clusters
$C_{\Psi}$ Capacity Score of LLM $\Psi$
$p_{\Psi}$ or $p$ Perception Score of LLM $\Psi$
Appendix B Limitation and Future Opportunities
Future Opportunities
Figure 8: COAT combines both strengths of LLMs that learn the rich knowledge of the world, and causal discovery methods to uncover the hidden causal world. With the uncovered causal knowledge, COAT can empower broader applications of causal methods.
Despite the progress in the past decades, existing causal discovery algorithms mainly rely on high-quality measured variables given by human experts [3, 4, 5]. However, the causal variables and their measurements are usually available in a wide range of real-world applications. For example, Amazon sellers who want to analyze the factors related to user ratings only have user reviews, generated by the underlying user preferences for certain product characteristics. Therefore, the lack of measured high-quality causal variables has been a major impediment to broader real-world applications of causal or causality-inspired methods [6].
This work establishes and demonstrates a preliminary implementation of the system Causal representatiOn AssistanT (COAT). We present comprehensive evaluations of COAT and find plentiful evidence that the recent emergence of Large Language Models (LLMs) [7, 8, 9, 10] has a great potential to mitigate the gap. In particular, as shown in Fig. 8, we envision an entangled system towards causal foundation models, which consists of two mutually beneficial components: LLMs and causal discovery methods (CDs). On the one hand, LLMs that learn rich world knowledge about the world can assist with discovering high-level hidden variables from low-level observational data [12], or even certain commonsense causal knowledge [13, 11, 14]. On the other hand, it has also been found that the reliability of LLMs in the reasoning of causality remains a debate [15, 16, 17, 13], due to a series of drawbacks of LLMs [18, 19, 20], as well as the ethical considerations [45, 46]. Nevertheless, CDs can uncover the underlying causal relations between the discovered high-level variables with guarantees. More importantly, CDs can also provide feedback to improve the identification of the variables. The combination of LLMs and CDs iteratively improves the discovery of the hidden causal world, and hence opens up a broader adoption of various causal methods.
Limitations
We would like to discuss some technical limitations of the current version of COAT.
  • Multicollinearity. Sometimes, LLM may output factors with shared similar semantic meanings or overlapped factor-value criteria. The consequence of this phenomenon is that multicollinearity could occur in the resulting structured data, or the data matrix (each sample per row) would not be full column rank. This would hinder the conditional independent tests. In experiments, we adopt numerical methods like PCA to drop similar factors. More sophisticated methods could be investigated in the future.
  • Toward complete causal graph. In this work, we mainly focus on one single target variable and identify a set of factors to serve as a Markov blanket of this target variable. Indeed, one can explore a more complete causal graph by applying COAT procedures on those identified factors to acquire a more comprehensive causal graph. One may also introduce multiple target variables or let LLM define a suitable target variable (and also parse it out) from pure raw observations. These extensions are also important, and we leave it to future work.
Meanwhile, identifiable causal discovery also requires certain assumptions about the data:
  • Faithfulness. The empirical distribution of the data reflects the actual data-generating process.
  • No selection bias. Otherwise, the condition of faithfulness would be violated.
  • Enough sample size. Our method involves statistical tests; the more, the better.
Appendix C Related Work
Causal discovery aims to discover the unknown causal relations from the observational data [4, 3], which is critical to both real-world applications and scientific discoveries [47, 23]. We use FCI [27] in this paper just to illustrate the idea. To attain identifiability better than the Markov equivalent class, one can choose more advanced Causal Discovery methods under different assumptions, like constrained functional [48, 49, 50], multiple domain data [51, 52, 53, 54, 55], and non-stationary data [56, 57, 51, 58]. There are also some works aiming to address realistic challenges, like structure models with arbitrary distribution [59], subsampled time series [60], and latent hierarchical causal structure  [61, 62, 63, 64]. Despite recent theoretical and empirical improvements [28, 5], most existing causal discovery approaches rely on well-structured data with human-crafted factors as inputs. They can easily suffer from the low quality of annotated data (e.g., latent confounders).
Causal representation learning aims to establish provable identifiability for latent high-level variables (like location and color of an object) from low-level observations such as images [6]. Such identification can be feasible with certain conditions, like auxiliary information [65, 66, 67], sparsity [68, 69, 70], or interventional data [71, 72, 73, 74, 75]. Recent works also generalize the causal representation learning to graph-structured data [76, 77, 78], and discuss the optimization challenges in realizing causal representation learning [79, 80]. In this work, we show that incorporating LLMs that learn world knowledge from massive training data could effectively relieve the need for artificial causal factor annotation. Meanwhile, COAT also opens up a new line to learn causal representations with rich pre-trained knowledge as well as feedback from causal discovery algorithms.
Reasoning with LLMs has achieved remarkable performance across a variety of tasks with few demonstrations of the samples [7, 8, 81, 10]. The strong capabilities of LLMs show that it is evident that LLMs could acquire and understand commonsense knowledge about the world [12]. The power of LLMs can be further unlocked with suitable context as inputs [29]. Nevertheless, LLMs have also been found to make mistakes in basic algorithmic reasoning [82, 83], easy to hallucinate nonfactual results [18], tend to learn shortcuts or dataset biases [84, 20, 19], vulnerable to adversarial jailbreaks or noisy interruptions [85, 86], and fall short in complex planning and reasoning tasks [12]. The drawbacks of LLMs render it risky to rely on the direct LLM reasoning results to derive any rigorous results. Therefore, we do not directly derive the results from LLMs. Rather, we merely leverage the learned world knowledge in LLMs to find useful causal factors by constructing proper instructions based on causal feedback.
Text mining of causal relations aims to analyze the causal relations implied by the semantics given a text [87, 88, 89]. In particular, some works focus on the identification of causal relations among events specified in documents [90, 91, 92]. Different from these tasks, where factors are entities or events described by text, in this work, COAT emphasizes crafting high-level factors that go beyond the unstructured data like text description and also establish identifiability on the Markov blanket for a given target variable. Note that the target variable is allowed to be not mentioned in the text or other unstructured data. Therefore, this work brings new scope and opportunities for these text-understanding tasks.
Causal learning with LLMs has received muchattention results from the community [11, 13]. Kiciman et al. [11] find that LLMs can recover the pairwise causal relations very well. Lampinen et al. [93] show that transformer-based agents can learn causal strategies passively if intervention is allowed during tests. Choi et al. [94], Long et al. [95], Ban et al. [96], Abdulaal et al. [14] propose to incorporate the causal discovery results by LLMs as a prior or constraint to improve the performance of data-driven causal discovery algorithms. However, Willig et al. [97], Zečević et al. [15] find that LLMs can not understand causality but simply retell the causal knowledge contained in the training data. Zhang et al. [13], Jin et al. [17, 16] find that LLMs can hardly provide satisfactory answers for discovering new knowledge or decision-making tasks. Although Long et al. [98] find that LLMs can build causal graphs with 3-4 nodes, Tu et al. [46] find that the performance of LLMs in more complex causal discovery remains limited as LLMs can hardly understand new concepts and knowledge. The aforementioned debate implies the limitations in directly adopting the causal discovery results by LLMs, which motivates us to incorporate the existing causal discovery algorithms with rigorous guarantees instead of LLMs to learn the causal relations.
LYU et al. [99] find that it is crucial to incorporate prompts aligned with the underlying causal story for LLMs to do pairwise causal relation inference. Zhang et al. [100] propose to leverage LLMs to audit causal discovery results in an explainable way.
The closest works to ours are Choi et al. [94], Long et al. [95], Ban et al. [96], Abdulaal et al. [14] which also incorporates LLMs into the pipeline of causal discovery. Nevertheless, all of the existing combinations of LLMs and causal discovery still focus on artificially curated structured data and rely on the capability of LLMs to infer causal relations, therefore, limited in both the reliability and the utility of LLMs in causal learning.
Appendix D Proofs for Theoretical Results
D.1 Proof for Proposition 2.2
Proposition D.1 (Restatement of Proposition 2.2).
Given the current representation as $h_{[k]}(X)=\left({\bm{w}}_{1}\left(X\right),{\bm{w}}_{2}\left(X\right),\cdots,{\bm{w}}_{k}\left(X\right)\right)$, and a new factor ${\bm{w}}_{k+1}$ satisfying
[…]
for Markov Blanket ${\mathcal{S}}\subseteq[k+1]$ of $Y$, i.e.,
$Y\perp\!\!\!\!\perp h_{[k+1]\setminus{\mathcal{S}}}(X)\mid h_{{\mathcal{S}}}(X),$ (13)
we have the following result about conditional mutual information:
$I\left(Y;X\mid h_{{\mathcal{S}}}(X)\right)=I\left(Y;X\mid h_{[k+1]}(X)\right)<I\left(Y;X\mid h_{[k]}(X)\right)$ (14)
Proof for Prop. 2.2.
From Eq. 12, we have
$H(y\mid h_{[k]}({\bm{x}});{\bm{w}}_{k+1}({\bm{x}}))<H(y\mid h_{[k]}({\bm{x}}));$ (15)
From Eq. 13, we have
$H(y\mid h_{{\mathcal{S}}}({\bm{x}});h_{[k+1]\setminus{\mathcal{S}}}({\bm{x}}))=H(y\mid h_{{\mathcal{S}}}({\bm{x}})).$ (16)
Therefore:
$H(y\mid h_{{\mathcal{S}}}({\bm{x}}))$ $=H(y\mid h_{{\mathcal{S}}}({\bm{x}}),h_{[k+1]\setminus{\mathcal{S}}}({\bm{x}}))$ (17)
$=H(y\mid h_{[k+1]}({\bm{x}}))$
$=H(y\mid h_{[k]}({\bm{x}}),{\bm{w}}_{k+1}({\bm{x}}))$
$<H(y\mid h_{[k]}({\bm{x}}))$
Also note that
$H(y\mid h_{{\mathcal{S}}}({\bm{x}}),{\bm{x}})=H(y\mid{\bm{x}})=H(y\mid h_{[k]}({\bm{x}}),{\bm{x}}),$ (18)
therefore:
$H(y\mid h_{{\mathcal{S}}}({\bm{x}}))-H(y\mid h_{{\mathcal{S}}}({\bm{x}}),{\bm{x}})<H(y\mid h_{[k]}({\bm{x}}))-H(y\mid h_{[k]}({\bm{x}}),{\bm{x}}),$ (19)
which is Eq. 14. ∎
D.2 Proof for Proposition 2.4
Proposition D.2 (Restatement of Proposition 2.4).
Given Assumption 2.3, for any small number $\epsilon,\delta\in(0,\frac{1}{2})$, with sufficiently $t$ rounds of COAT, i.e.,
$\sqrt{t}>\frac{|z_{\delta}|\sqrt{1-p}}{2\sqrt{p}}\left(1+\sqrt{1+\frac{4\log{\epsilon}}{z_{\delta}^{2}(1-p)\log{1-C_{\Psi}}}}\,\right),$ (20)
where $z_{\delta}$ is the $\delta$-quantile of the standard normal distribution, we have
$\Pr\left(\frac{I\left(Y;X\mid h_{\leq t}(X)\right)}{I\left(Y;X\right)}<\epsilon\right)\geq 1-\delta.$ (21)
Proof for Prop. 2.4.
Let $n_{s}$ be the number of rounds that LLM proposed at least one usable factor satisfying Assumption 2.3. Since $t$ is large, its Binomial distribution $\text{Binom}(t,p)$ can be approximated by Gaussian distribution ${\mathcal{N}}\left(tp,tp(1-p)\right)$. To enforce
$\Pr\left(\frac{I\left(y;{\bm{x}}\mid h_{\leq t}({\bm{x}})\right)}{I\left(y;{\bm{x}}\right)}<\epsilon\right)$ $\geq\Pr\Big{(}\left(1-C_{\Psi}\right)^{n_{s}}<\epsilon\Big{)}$ (22)
$=\Pr\Big{(}n_{s}>\frac{\log{\epsilon}}{\log{(1-C_{\Psi})}}\Big{)}$
$=\Pr\Big{(}\frac{n_{s}-tp}{\sqrt{tp(1-p)}}>\frac{1}{\sqrt{tp(1-p)}}\left(\frac{\log{\epsilon}}{\log{(1-C_{\Psi})}}-tp\right)\Big{)}$
$\geq 1-\delta,$
we have
$\frac{1}{\sqrt{tp(1-p)}}\left(\frac{\log{\epsilon}}{\log{(1-C_{\Psi})}}-tp\right)<z_{\delta}=-|z_{\delta}|$ (23)
with Gaussian distribution approximation.
Isolating $\sqrt{t}>0$ from the above inequality, we would have the desired result. ∎
Appendix E More Details about Experiments
E.1 More Details on Constructing AppleGastronome
In the AppleGastronome benchmark, we consider the target variable as a rating score of the apple by several gastronomes. Each apple has its own attributes, including size, smell, and taste (or sweetness). Each gastronome has a unique preference for some attributes of the apple. They will give and rating as well as write a review according to the matchness of the apple with respect to their preference. We generate the review using GPT 4 by fetching GPT 4 the preferences and the apple attributes.
The prompts for generating the unstructured inputs are given in Fig. 9. The additional results on Relation Extraction are given in Table. 3.
Table 3: Causal relation extraction results in AppleGastronome.
LLM Method SHD SID Recall Precision F1
GPT 4o pairwise 5.33$\pm$2.62 0.00$\pm$0.00 1.00$\pm$0.00 0.45$\pm$0.07 0.62$\pm$0.07
COAT 1.00$\pm$0.00 0.00$\pm$0.00 1.00$\pm$0.00 0.79$\pm$0.03 0.89$\pm$0.02
GPT 4 pairwise 8.67$\pm$2.05 1.33$\pm$1.89 0.93$\pm$0.09 0.35$\pm$0.05 0.51$\pm$0.06
COAT 2.33$\pm$0.47 1.00$\pm$1.41 0.93$\pm$0.09 0.70$\pm$0.09 0.79$\pm$0.04
GPT 3.5 pairwise 5.67$\pm$4.03 3.00$\pm$2.94 0.67$\pm$0.24 0.49$\pm$0.36 0.55$\pm$0.32
COAT 2.00$\pm$0.82 2.67$\pm$2.05 0.75$\pm$0.20 0.63$\pm$0.21 0.68$\pm$0.21
Mistral-Large pairwise 7.33$\pm$1.89 0.00$\pm$0.00 1.00$\pm$0.00 0.38$\pm$0.03 0.55$\pm$0.03
COAT 2.00$\pm$0.00 1.33$\pm$1.25 0.85$\pm$0.11 0.74$\pm$0.05 0.78$\pm$0.02
Mistral-Medium pairwise 6.00$\pm$1.63 2.00$\pm$0.00 0.51$\pm$0.13 0.37$\pm$0.04 0.42$\pm$0.07
COAT 0.33$\pm$0.47 0.00$\pm$0.00 1.00$\pm$0.00 0.94$\pm$0.08 0.97$\pm$0.04
llama-3-70b pairwise 4.00$\pm$1.41 0.33$\pm$0.47 0.92$\pm$0.12 0.51$\pm$0.09 0.65$\pm$0.11
COAT 2.33$\pm$0.47 3.67$\pm$1.25 0.75$\pm$0.00 0.70$\pm$0.07 0.72$\pm$0.04
llama-2-70b pairwise 3.33$\pm$1.25 1.00$\pm$1.41 0.89$\pm$0.16 0.48$\pm$0.11 0.62$\pm$0.13
COAT 1.00$\pm$0.82 0.33$\pm$0.47 0.89$\pm$0.16 0.81$\pm$0.14 0.84$\pm$0.14
Qwen-1.5-110B pairwise 6.67$\pm$2.62 2.33$\pm$1.25 0.62$\pm$0.03 0.38$\pm$0.09 0.47$\pm$0.07
COAT 4.00$\pm$1.63 7.00$\pm$4.32 0.47$\pm$0.34 0.47$\pm$0.34 0.70$\pm$0.10
DeepSeek-V2 pairwise 5.00$\pm$1.41 1.33$\pm$1.89 0.93$\pm$0.09 0.47$\pm$0.05 0.62$\pm$0.06
COAT 1.33$\pm$1.25 2.33$\pm$1.70 0.85$\pm$0.11 0.87$\pm$0.19 0.85$\pm$0.14
Claude-3-Opus pairwise 7.33$\pm$1.25 0.33$\pm$0.47 0.93$\pm$0.09 0.40$\pm$0.05 0.56$\pm$0.07
COAT 1.33$\pm$0.47 0.00$\pm$0.00 1.00$\pm$0.00 0.79$\pm$0.06 0.88$\pm$0.04
Table 4: Independence tests of the annotation noises with annotated features and other noises AppleGastronome.
LLM Test Object T P-value
GPT 4 Feature 0.2828 0.9997
Noise 0.0327 0.9962
GPT 3.5 Feature 0.4803 0.0325
Noise 0.0446 0.9962
Figure 9: Illustration of prompts for generating AppleGastronome.
Examples of AppleGastronome are given in Fig. 10.
Figure 10: Illustration of examples in AppleGastronome.
E.2 More Details on Prompts for AppleGastronome
The prompts for factor proposal are given in Fig. 11.
Figure 11: Illustration of the prompt for factor proposal.
The prompt for factor annotation is given in Fig. 12.
Figure 12: Illustration of the prompt for factor annotation.
The prompts for constructing feedback are given in Fig. 13.
Figure 13: Illustration of the prompt for feedback.
E.3 More Details of Results on AppleGastronome
The detailed causal graph results are given from Fig. 14 to Fig. 18.
Independent tests about annotation on the Apple Gastronome benchmark are shown in Table  4.
E.4 The Full Result on the Apple Gastronome Benchmark
Full Results on the Apple Gastronome Benchmark are shown in Table 5.
Full Results of Causal Metrics each Round of each LLM on Apple Gastronome Benchmark are shown in Table 6.
Table 5: Full Results on the Apple Gastronome Benchmark.
LLMs Method MB NMB OT Recall Precition F1
GPT 4o META 3.00$\pm$0.82 1.00$\pm$0.00 4.00$\pm$0.82 0.60$\pm$0.16 0.37$\pm$0.09 0.46$\pm$0.12
DATA 3.00$\pm$0.00 0.67$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.83$\pm$0.12 0.69$\pm$0.04
DATA+CoT 4.67$\pm$0.58 1.00$\pm$0.00 0.33$\pm$0.58 0.93$\pm$0.12 0.78$\pm$0.10 0.85$\pm$0.10
COAT 4.00$\pm$0.82 0.00$\pm$0.00 0.00$\pm$0.00 0.80$\pm$0.16 1.00$\pm$0.00 0.88$\pm$0.10
GPT 4 META 2.67$\pm$0.94 0.67$\pm$0.47 2.33$\pm$0.47 0.53$\pm$0.19 0.46$\pm$0.08 0.49$\pm$0.13
DATA 3.00$\pm$0.00 0.33$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.92$\pm$0.12 0.72$\pm$0.04
DATA+CoT 4.33$\pm$0.58 0.83$\pm$0.29 0.17$\pm$0.29 0.87$\pm$0.12 0.81$\pm$0.02 0.84$\pm$0.06
COAT 4.00$\pm$0.82 0.33$\pm$0.47 0.00$\pm$0.00 0.80$\pm$0.16 0.93$\pm$0.09 0.85$\pm$0.11
GPT 3.5 META 3.33$\pm$1.25 0.33$\pm$0.47 4.33$\pm$1.25 0.67$\pm$0.25 0.42$\pm$0.12 0.51$\pm$0.15
DATA 2.67$\pm$0.47 0.67$\pm$0.47 0.00$\pm$0.00 0.53$\pm$0.09 0.81$\pm$0.14 0.64$\pm$0.10
DATA+CoT 5.00$\pm$0.00 1.00$\pm$0.00 1.33$\pm$0.58 1.00$\pm$0.00 0.68$\pm$0.05 0.81$\pm$0.04
COAT 3.67$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.73$\pm$0.09 1.00$\pm$0.00 0.84$\pm$0.07
Mistral-Large META 2.00$\pm$0.82 0.67$\pm$0.47 2.67$\pm$0.94 0.40$\pm$0.16 0.37$\pm$0.09 0.38$\pm$0.12
DATA 3.00$\pm$0.00 0.33$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.92$\pm$0.12 0.72$\pm$0.04
DATA+CoT 4.33$\pm$0.58 1.00$\pm$0.00 0.67$\pm$0.58 0.87$\pm$0.12 0.73$\pm$0.07 0.79$\pm$0.05
COAT 4.33$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.87$\pm$0.09 1.00$\pm$0.00 0.93$\pm$0.05
Mistral-Medium META 3.00$\pm$0.00 0.67$\pm$0.47 1.67$\pm$1.25 0.60$\pm$0.00 0.59$\pm$0.13 0.59$\pm$0.07
DATA 3.00$\pm$0.00 0.67$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.83$\pm$0.12 0.69$\pm$0.04
DATA+CoT 4.33$\pm$0.58 1.00$\pm$0.00 0.67$\pm$0.58 0.87$\pm$0.12 0.73$\pm$0.07 0.79$\pm$0.05
COAT 4.67$\pm$0.47 0.00$\pm$0.00 0.00$\pm$0.00 0.93$\pm$0.09 1.00$\pm$0.00 0.96$\pm$0.05
llama-3-70b META 2.67$\pm$0.47 0.33$\pm$0.47 4.67$\pm$0.47 0.53$\pm$0.09 0.35$\pm$0.06 0.42$\pm$0.07
DATA 2.67$\pm$1.25 0.33$\pm$0.47 0.00$\pm$0.00 0.53$\pm$0.25 0.93$\pm$0.09 0.63$\pm$0.21
DATA+CoT 2.67$\pm$0.58 0.67$\pm$0.58 1.00$\pm$0.00 0.53$\pm$0.12 0.62$\pm$0.13 0.57$\pm$0.11
COAT 3.67$\pm$0.47 0.33$\pm$0.47 0.00$\pm$0.00 0.73$\pm$0.09 0.93$\pm$0.09 0.81$\pm$0.06
llama-2-70b META 2.33$\pm$0.47 0.67$\pm$0.47 4.67$\pm$1.25 0.47$\pm$0.09 0.32$\pm$0.07 0.37$\pm$0.06
DATA 2.33$\pm$0.94 0.67$\pm$0.47 0.00$\pm$0.00 0.47$\pm$0.19 0.75$\pm$0.20 0.57$\pm$0.20
DATA+CoT 3.00$\pm$1.73 0.67$\pm$0.58 0.33$\pm$0.58 0.60$\pm$0.35 0.71$\pm$0.34 0.65$\pm$0.35
COAT 3.00$\pm$0.00 0.67$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.83$\pm$0.12 0.69$\pm$0.04
Qwen-1.5-110B META 2.00$\pm$0.00 1.00$\pm$0.00 4.00$\pm$0.00 0.40$\pm$0.00 0.29$\pm$0.00 0.33$\pm$0.00
DATA 3.00$\pm$0.82 1.00$\pm$0.00 0.00$\pm$0.00 0.60$\pm$0.16 0.74$\pm$0.05 0.66$\pm$0.12
DATA+CoT 3.67$\pm$1.53 0.67$\pm$0.58 0.00$\pm$0.00 0.73$\pm$0.31 0.83$\pm$0.17 0.77$\pm$0.23
COAT 4.00$\pm$0.82 0.33$\pm$0.47 0.00$\pm$0.00 0.80$\pm$0.16 0.93$\pm$0.09 0.85$\pm$0.11
DeepSeek-V2 META 2.33$\pm$0.47 1.00$\pm$0.00 3.00$\pm$0.82 0.47$\pm$0.09 0.37$\pm$0.03 0.41$\pm$0.04
DATA 3.33$\pm$0.47 0.67$\pm$0.47 0.00$\pm$0.00 0.67$\pm$0.09 0.85$\pm$0.11 0.74$\pm$0.05
DATA+CoT1 5.00$\pm$0.00 1.00$\pm$0.00 0.67$\pm$0.58 1.00$\pm$0.00 0.75$\pm$0.07 0.86$\pm$0.04
COAT 3.67$\pm$0.47 0.67$\pm$0.47 0.00$\pm$0.00 0.73$\pm$0.09 0.87$\pm$0.09 0.78$\pm$0.02
Claude-3-Opus META 2.00$\pm$0.00 1.00$\pm$0.00 3.00$\pm$1.41 0.40$\pm$0.00 0.35$\pm$0.07 0.37$\pm$0.04
DATA 3.33$\pm$0.47 0.33$\pm$0.47 0.00$\pm$0.00 0.67$\pm$0.09 0.93$\pm$0.09 0.77$\pm$0.02
DATA+CoT 2.67$\pm$0.58 0.33$\pm$0.58 1.00$\pm$1.00 0.53$\pm$0.12 0.67$\pm$0.14 0.59$\pm$0.13
COAT 5.00$\pm$0.00 0.33$\pm$0.47 0.00$\pm$0.00 1.00$\pm$0.00 0.94$\pm$0.08 0.97$\pm$0.04
1DeepSeek-V2 is no more available when we add this baseline, therefore, we use DeepSeek-V2.5 instead.
Table 6: Full Result of Causal Metrics each Round each LLM on Apple Gastronome Benchmark
LLMs COAT Round Perception Score Capacity Score $I(y;x|h_{S})$
GPT 4o iter 1 0.82$\pm$0.02 0.24$\pm$0.01 0.28$\pm$0.02
iter 2 0.44$\pm$0.08 0.26$\pm$0.02 0.21$\pm$0.03
iter 3 0.38$\pm$0.12 0.19$\pm$0.09 0.24$\pm$0.05
GPT 4 iter 1 0.78$\pm$0.02 0.18$\pm$0.02 0.45$\pm$0.04
iter 2 0.61$\pm$0.08 0.26$\pm$0.05 0.28$\pm$0.05
iter 3 0.39$\pm$0.28 0.22$\pm$0.00 0.29$\pm$0.02
GPT 3.5 iter 1 0.81$\pm$0.02 0.21$\pm$0.01 0.42$\pm$0.05
iter 2 0.52$\pm$0.11 0.23$\pm$0.02 0.31$\pm$0.04
iter 3 0.38$\pm$0.10 0.21$\pm$0.03 0.33$\pm$0.11
Mistral-Large iter 1 0.78$\pm$0.02 0.20$\pm$0.01 0.41$\pm$0.04
iter 2 0.61$\pm$0.08 0.20$\pm$0.03 0.29$\pm$0.02
iter 3 0.33$\pm$0.33 0.32$\pm$0.00 0.25$\pm$0.05
Mistral-Medium iter 1 0.78$\pm$0.02 0.20$\pm$0.02 0.39$\pm$0.05
iter 2 0.39$\pm$0.15 0.19$\pm$0.10 0.32$\pm$0.04
iter 3 0.36$\pm$0.10 0.37$\pm$0.14 0.25$\pm$0.08
llama-3-70b iter 1 0.75$\pm$0.00 0.19$\pm$0.02 0.47$\pm$0.08
iter 2 0.36$\pm$0.31 0.19$\pm$0.00 0.38$\pm$0.02
iter 3 0.27$\pm$0.38 0.18$\pm$0.00 0.36$\pm$0.03
llama-2-70b iter 1 0.77$\pm$0.02 0.21$\pm$0.05 0.43$\pm$0.13
iter 2 0.08$\pm$0.12 0.35$\pm$0.00 0.35$\pm$0.03
iter 3 0.28$\pm$0.04 0.14$\pm$0.07 0.33$\pm$0.03
Qwen-1.5-110B iter 1 0.77$\pm$0.02 0.17$\pm$0.01 0.45$\pm$0.09
iter 2 0.56$\pm$0.16 0.21$\pm$0.12 0.32$\pm$0.08
iter 3 0.28$\pm$0.21 0.18$\pm$0.02 0.33$\pm$0.11
DeepSeek-V2 iter 1 0.89$\pm$0.08 0.17$\pm$0.02 0.42$\pm$0.05
iter 2 0.62$\pm$0.10 0.20$\pm$0.02 0.34$\pm$0.02
iter 3 0.67$\pm$0.24 0.40$\pm$0.12 0.29$\pm$0.06
Claude-3-Opus iter 1 0.77$\pm$0.02 0.18$\pm$0.00 0.43$\pm$0.06
iter 2 0.56$\pm$0.08 0.39$\pm$0.11 0.24$\pm$0.06
iter 3 0.75$\pm$0.25 0.15$\pm$0.00 0.22$\pm$0.05
E.5 Implementation of the FCI algorithm
We use a third-party open-sourced Python library to perform the FCI algorithm: https://causal-learn.readthedocs.io/en/latest/
We set $\alpha=0.05$ , and independence_test_method="fisherz" hroughout all experiments. Other parameters are kept as the default.
(a) Ground truth
(b) FCI results
Figure 14: Ground truth and faithful (via FCI algorithm) causal graphs in AppleGastronome.
(a) GPT 4 reasoning
(b) GPT 4 COAT
Figure 15: Causal graphs with GPT 4 in AppleGastronome.
(a) GPT 3.5 reasoning
(b) GPT 3.5 COAT
Figure 16: Causal graphs with GPT 3.5 in AppleGastronome.
(a) LLaMA2 reasoning
(b) LLaMA2 COAT
Figure 17: Causal graphs with Llama-2 in AppleGastronome.
(a) Mistral reasoning
(b) Mistral COAT
Figure 18: Causal graphs with mistral Medium in AppleGastronome.
(a) Claude-3-Opus reasoning
(b) Claude-3-Opus COAT
Figure 19: Causal graphs with Claude-3-Opus in AppleGastronome.
(a) DeepSeek-V2 reasoning
(b) DeepSeek-V2 COAT
Figure 20: Causal graphs with DeepSeek-V2 in AppleGastronome.
(a) Llama-3-70b reasoning
(b) Llama-3-70b COAT
Figure 21: Causal graphs with Llama-3-70b in AppleGastronome.
(a) mistral-Large reasoning
(b) mistral-Large COAT
Figure 22: Causal graphs with mistral-Large in AppleGastronome.
(a) qwen-1.5-110B reasoning
(b) qwen-1.5-110B COAT
Figure 23: Causal graphs with qwen-1.5-110B in AppleGastronome.
(a) GPT-4o reasoning
(b) GPT-4o COAT
Figure 24: Causal graphs with GPT-4o in AppleGastronome.
E.6 More Details on Constructing Neuropathic
Figure 25: Illustration of prompts for generating Neuropathic.
In the Neuropathic benchmark, we convert the dataset into a clinical diagnosis task. In the original dataset, there are three levels of causal variables, including the symptom-level, radiculopathy-level and the pathophysiology-level. In experiments, we mainly consider the target variable of right shoulder impingement. When generating the clinical diagnosis notes as ${\bm{x}}$ using GPT 4, we will avoid any mentioning of variables other than symptoms.
As we intend to leverage the Neuropathic benchmark to simulate the real-life diagnosis, after the factor proposal stage, we directly incorporate external experts that measure the values of the candidate factors. The prompts to generate the diagnosis records are given in Fig. 25.
Examples of Neuropathic are given in Fig. 26.
Figure 26: Illustration of examples in Neuropathic.
E.7 More Details of Results on Neuropathic
The detailed causal graph results are given from Fig. 27 to Fig. 31.
(a) Ground truth
(b) FCI results
Figure 27: Ground truth and faithful (via FCI algorithm) causal graphs in Neuropathic.
(a) GPT 4 reasoning
(b) GPT 4 COAT
Figure 28: Causal graphs with GPT 4 in Neuropathic.
(a) GPT 3.5 reasoning
(b) GPT 3.5 COAT
Figure 29: Causal graphs with GPT 3.5 in Neuropathic.
(a) LLaMA-2-70b reasoning
(b) LLaMA-2-70b COAT
Figure 30: Causal graphs with LLaMA-2-70b in Neuropathic.
(a) Mistral-Medium reasoning
(b) Mistral-Medium COAT
Figure 31: Causal graphs with Mistral-Medium in Neuropathic.
E.8 Discussion on the time complexity
Assume there are m samples with n possible factors.
For Factor Proposal:
META only needs to interact once with LLM, so it is O(1). DATA runs one single COAT round, so it is O(1). COAT interacts with LLMs multiple rounds. In each round, at least one new factor should be proposed; otherwise, the loop will stop. So it is O(n) For Factor Annotation:
META: Not applicable. DATA: At most n factors would be proposed in a single round. And each of them needs m times annotations by LLMs for all samples. So it is O(nm). COAT: At most n factors would be proposed during all rounds. And each of them needs m times annotations by LLMs for all samples. So it is O(nm). For Causal Discovery:
Pair-wise reasoning by LLMs: O($n^{2}$) COAT: LLM is not involved. The computational cost of it depends on the numeric methods. The FCI algorithm used by COAT has a time complexity that goes exponentially with n. Learning the causal graph over a large number of nodes effectively is still an open problem in causal discovery literature.
E.9 Resources
We utilized a system comprising two Intel Xeon E5-2630v4 processors with 2.2GHz, two NVIDIA Tesla P40 GPUs, and 256 GB of memory. For conversations with large language models (LLMs), we leveraged the poe.com platform, while annotations were facilitated using the OpenAI API and the Mistral API.
Appendix F COAT with Different Causal Discovery Algorithm
Table 7: Causal discovery results in Neuropathic. PA, AN, and OT refer to the parents, ancestors, and others, respectively. Accuracy and F1 measure the recovery of the causal ancestors.
LLM Method Factor Proposal
PA AN OT Acc F1
GPT 4 Meta 3 5 6 0.91 0.59
Data 2 2 0 0.95 0.50
COAT 3 6 3 0.96 0.80
Data (LiNGAM) 3 3 0 0.96 0.67
COAT (LiNGAM) 3 6 4 0.95 0.75
GPT 3.5 Meta 3 5 6 0.91 0.59
Data 3 5 4 0.94 0.67
COAT 3 5 2 0.96 0.77
Data (LiNGAM) 2 3 4 0.91 0.46
COAT (LiNGAM) 3 5 4 0.94 0.67
LLaMA2-70b Meta 2 4 5 0.91 0.53
Data 3 3 1 0.95 0.60
COAT 3 6 2 0.97 0.86
Data (LiNGAM) 3 4 4 0.92 0.57
COAT (LiNGAM) 3 6 4 0.95 0.75
Mistral-Med Meta 3 6 3 0.96 0.80
Data 3 3 2 0.94 0.66
COAT 3 6 2 0.97 0.86
Data (LiNGAM) 3 3 3 0.92 0.50
COAT (LiNGAM) 3 6 3 0.96 0.80
(a) Faithful ground truth
(b) GPT-4 COAT with LiNGAM
(c) GPT-3.5 COAT with LiNGAM
(d) llama-2-70b COAT with LiNGAM
(e) Mistral-med COAT with LiNGAM
Figure 32: The discovered causal graphs in Neuropathic with LiNGAM.
Appendix G Ablation Study
Group Size in Prompt. In the COAT prompt, several samples are given and grouped by the values of the target variable. The samples in each group are randomly selected to a fixed number (like 3 samples per group). Empirically, we keep it to be 3 throughout all experiments (sometimes smaller than 3 if samples are not enough). In practice, it is mainly constrained by the LLM’s context length.
The Number of Clusters. When constructing feedback, we first use clustering to separate the dataset and then find the cluster where the target variable is not explained well by current factors (This is a heuristic for the problem in line 191). Empirically, we set the number of clusters to be one plus the number of current factors.
We conduct ablation studies of COAT with GPT-4 using different hyperparameters:
Table 8: Results about Ablation Study on Hyperparameters.
Method Cluster Size Group Size MB NMB OT Recall Precision F1
META - - 2.67$\pm$0.94 0.67$\pm$0.47 2.33$\pm$0.47 0.53$\pm$0.19 0.46$\pm$0.08 0.49$\pm$0.13
DATA - 3 3.00$\pm$0.00 0.33$\pm$0.47 0.00$\pm$0.00 0.60$\pm$0.00 0.92$\pm$0.12 0.72$\pm$0.04
COAT len(factor)+1 3 4.00$\pm$0.82 0.33$\pm$0.47 0.00$\pm$0.00 0.80$\pm$0.16 0.93$\pm$0.09 0.85$\pm$0.11
COAT len(factor)+1 1 4.67$\pm$0.58 0.00$\pm$0.00 0.00$\pm$0.00 0.93$\pm$0.12 1.00$\pm$0.00 0.96$\pm$0.06
COAT 2 3 3.67$\pm$1.53 0.00$\pm$0.00 0.00$\pm$0.00 0.73$\pm$0.31 1.00$\pm$0.00 0.82$\pm$0.22
As shown in Table 8, one can observe that COAT is not sensitive to these hyperparameters and performs robustly well than the baselines under different hyperparameter setups.
Prompt Template. We conduct an ablation study with a different prompt template following [101]:
  • We put the task description (also the format instructions) in the beginning [System] part, and we put samples in the last [Data] part of the prompt.
  • The markdown grammar is replaced by blankets to represent headings, like [System] , [Data] , and [Groups with Y=1]
  • 3 COAT iterations are performed, which is aligned with the original experimental setup.
Table 9: COAT with changed prompt template
Model MB NMB OT Recall Precision F1
GPT-4 4 0 0 0.80 1.00 0.89
GPT-3.5-Turbo 4 0 0 0.80 1.00 0.89
Mistral-Medium 3 0 0 0.60 1.00 0.75
In Table 9, we observe that COAT is robust to the choice of templates, rejects unexpected factors (zero NMB and OT), and keeps a high precision.
Appendix H Summary of Benchmark Data
Benchmark Name Type Source Sample Sample Size Ground Truth
Apple Gastronome Synthetic Data This paper. It will be open-sourced under CC-BY 4.0. Textual 200 textual reviews. Yes
Neuropathic Semi Real-world Data Tabular samples from Tu et al. [26] on their GitHub repo under CC-BY 4.0. Textual samples are generated by this paper. Textual 100 synthetic textual data for prompt construction; 1000 tabular data for CI tests. Yes
Brain Tumor Real-world Data An open Kaggle dataset (kaggle/brain-tumor-classification-mri) with an open-sourced project. [38]. MIT License. Image 60 MRI samples for COAT and CI tests. The sample size is small due to the expensive API cost. One can easily construct a much larger dataset by following our instructions. No
Stock News Real-world Data An open Kaggle dataset (kaggle/stock-price-and-news-realted-to-it[39]. MIT License. Textual with time index 200 samples for COAT experiments. 400 samples for empirical validation. No
ENSO Real-world Data The NOAA 20th Century Reanalysis V3 dataset [40] from their website at https://psl.noaa.gov (with CC0 1.0 License). We developed a set of simple tools for LLMs to utilize the official database. NetCDF Monthly observation for 1836/01 to 2015/12; 1.0 degree latitude x 1.0 degree longitude global grid (360x181); 90+ climate variables. No
Table 10: Summary of Benchmark Data
Appendix I Case Study on Brain Tumor
In this section, we utilize the COAT to explore the image dataset.
I.1 The Brain Tumor Dataset
Dataset
Magnetic Resonance Imaging (MRI) is an important technique for detecting tumors in the human brain. The images are from an open Kaggle dataset (kaggle/brain-tumor-classification-mri) with an open-sourced project. [38]. Each sample is a scanning MRI of a human brain. In this case study, the interesting variable is the tumor type. We consider three types of MRI images: glioma, meningioma, and no tumor. We include 20 images for each category and the total sample size is 60.
Figure 33: The initial input to COAT in the Brain Tumor case study. Each row contains 5 samples randomly selected from one category (top-down: glioma, meningioma, and no tumor).
Data Processing
We use gpt-4-vision-preview to handle image samples. As the current gpt4 cannot process multiple images simultaneously, we concatenate samples from different categories into one picture, as shown in Fig. 33, and give instructions in the prompt to explain the format, as shown in Fig. 34. For each proposed factor, the LLMs would go through all 60 sample images individually to evaluate the factor value.
Figure 34: Illustration of the prompt for the Brain Tumor case study. It contains instructions about understanding the combined input picture.
Evaluation
The ground truth is not directly available to this dataset. Therefore, each factor in the final result will be evaluated by searching the related literature and will be explained by human interpretation.
I.2 Result and Discussion
Result interpretation
As shown in Fig. 35, there are two visual factors that appeared in the final causal graph: contrast enhancement and mass effect. The first factor (contrast enhancement) is about whether the bright part is uniform or heterogeneous in the sample image. Another factor (mass effect) is about whether the tumor tissue has caused significant displacement of normal brain structures. One may refer to Fig. 36 to see detailed descriptions of these two factors.
Figure 35: Final causal graph by COAT in the Brain Tumor case study.
Figure 36: Detailed descriptions of these two final factors in the Brain Tumor case study. These descriptions were also directly fed to LLMs for factor annotation on images.
Figure 37: Medical literature about the brain tumor. Searched according to keywords in factor descriptions.
Justification for the final factors
We verify the proposed factors by the medical literature  [102, 103, 104, 105], as shown Fig. 37. Note that these papers are searched according to keywords in factor descriptions and, therefore, can exclude some relevant studies. Therefore, the conclusions of this preliminary exploration should be considered a reference point for further in-depth validation by domain experts.
The factor description for contrast enhancement directly matches with the related paper (the first two papers displayed in Fig. 37). In addition, this can be visually checked in Fig. 33. Therefore, we believe it is a good factor.
The second factor, mass effect, may not directly match the papers. It is pointed out in the last two papers displayed in Fig. 37 that the two tumors have different axial locations and thus influence the magnitude of the displacement of tissues. Both the keywords ”axial location” and ”displacement” have occurred in its factor description.
Appendix J Case Study on Stock News
In this section, we utilize COAT to explore time series data with text.
J.1 The Stock News Dataset
Dataset
The dataset consists of the stock value of Microsoft (MSFT) from 2006 to 2009 and its news summary (only include news in the New York Times). This is a subset of an open Kaggle dataset (kaggle/stock-price-and-news-realted-to-it). Each sample is one trading day with the company’s close stock price and news. The target value is the future return rate, and we are curious about factors in the related news. We fed data during the first 200 trading days to COAT, and we used the following 400 trading days for evaluation.
Data processing
The future return rate is calculated by close prices:
$r_{t}=\frac{\text{close}_{t+4}}{\text{close}_{t}}-1.$ (24)
The target value is a binary variable according to the return rate
$Y_{t}=\mathbf{1}_{r_{t}>0},$ (25)
where $\mathbf{1}_{A}$ is the indicator function. News fed to COAT are grouped according to the value of $Y_{t}$. To keep the whole prompt within the context limit, only 3 samples would be randomly included in each group.
Factor processing
Given one proposed factor, denote its annotated value at the $t$-th trading day as $a_{t}\in\{-1,0,1\}$. Each factor is rolling averaged over the past $M$ days:
$S_{t}=\frac{1}{M}\sum_{t-M\geq i\geq t}a_{i}.$ (26)
Each variable, including return rate, is normalized by the rolling standard deviation over the past $M$ days:
$F_{t}=\frac{1}{\text{Sd}\left(\left\{S_{i}\right\}_{t-M\geq i\geq t}\right)+1}S_{t}.$ (27)
An example of a processed factor can be seen in Fig. 38.
Figure 38: Example of one processed factors Innovation and Technology Focus during the first 200 trading days. The red star marks the highest value. In June 2007, there was a discussion about future competition in desktop operating systems and the trend towards web-based services [106].
J.2 Result and Discussion
Result Causal Graph
As shown in Fig. 39, four factors are identified in the final causal graph to form a Markov blanket of return rate: Product Focus, Legal and Regulatory Issues, Market Strategy, and Innovation and Technology Focus. One can refer to Fig. 40 for more detailed descriptions of these factors. One factor (Innovation and technology focus) is identified as a possible cause of the return rate; this matches the nature of the company’s type and reflects people’s expectation of the company to keep creating innovative computer software. It is also interesting to see that COAT captures the structure between factors and market strategy, where product focus and legal and regulatory issues are identified as potential causes. It also implies the existence of a latent confounder between market strategy and return rate that may not be significantly reflected in news text.
Figure 39: The final causal graph by COAT in the Stock News case study
Figure 40: The detailed factor descriptions proposed by COAT in the Stock News case study
Evaluation by Trading Strategies
For each factor $\{F_{t}\}$ processed in Eq. 27, we establish a trading strategy on it to see the performance in the out-of-sample trading days. At each trading day $t$, we fit the following model
$r_{t}=\alpha+\beta F_{t},$ (28)
using samples $\{r_{i},F_{i}\}_{t-200\leq i<t}$. Note that $r_{t}$ is the future return rate defined in Eq. 24 and is not available on this day. After the estimation, we make decision based on $\hat{r}_{t}:=\hat{\alpha}+\hat{\beta}F_{t}$. If $\hat{r}_{t}>0$, we go long with 1 unit capital, i.e., purchase that amount of stock, the gain would be gain would be $r\times 1$ units capital. If $\hat{r}_{t}<0$, we go short with 1 unit capital, i.e., borrow and sell that amount of stocks immediately and buy them back next time, the gain would be $-r\times 1$ units capital.
To align with the definition in Eq. 24, we make trading decisions every $4$ trading day. We introduce an additional Buy and Hold baseline to always go long one unit capital. The trading evaluation is after the $200$-th day and thus has no overlapping with data fed to COAT.
The cumulative return plot is shown in Fig. 41, and the metrics commonly used in economic literature [107, 108, 109] are shown in Table 11. One important metric to see a trading strategy’s effectiveness is the sharp ratio: a measure of risk-adjusted return, showing the excess return (over the risk-free rate, we set it to be $2\%$) per unit of standard deviation. A higher Sharpe ratio indicates better performance per unit of risk. We see that the Innovation and Technology Focus factor yields the highest sharp ratio and outperforms other non-causal factors.
Figure 41: The out-of-sample cumulative return of trading strategies based on different factors in the Stock News case study
Table 11: Performance about trading strategy according to each factors
Buy and Hold product focus legal/regulatory issues market strategy innovation and technology focus
Expected Return -3.50 0.29 0.38 0.48 0.72
Sharp Ratio -2.38 0.70 0.94 1.20 1.86
t-Stat -3.00 0.89 1.19 1.51 2.35
Information Ratio - 2.45 2.50 2.56 2.73
$\alpha$ 3.79 3.88 3.98 4.22
$\alpha$ t-Stat - 3.09 3.15 3.23 3.44
Max Loss -0.50 -0.15 -0.15 -0.15 -0.15
Skew -0.43 -0.14 -0.15 -0.41 -0.38
Appendix K Case Study on El Niño–Southern Oscillation (ENSO)
In this section, we use COAT to explore a scientific problem using open datasets in climate science. The case study shows its potential to handle more complex data formats when combined with tools.
The ENSO Background
El Niño-Southern Oscillation (ENSO) is an important phenomenon in the Pacific Ocean and is characterized by irregular fluctuations in sea surface temperatures[110]. ENSO consists of the warm phase (known as El Niño) and the cold phase (known as La Niña). It profoundly influences global weather patterns, including precipitation, storm development, and temperature anomalies [111]. The prediction of ENSO events involves the complex interplay between oceanic and atmospheric systems and, therefore, is still an open problem [112, 113].
NOAA Dataset
The NOAA 20th Century Reanalysis V3 dataset contains high-dimensional information about Earth’s atmosphere with fine-grained time and space coverage from the 19th century to the early 21st century [40]. This dataset amalgamates many observational data and uses the ensemble filter data assimilation method to reconstruct the historical state of the global atmosphere. The dataset includes a fine-grained spatial coverage with $360\times 181$ grids. We only use monthly data in this case study.
K.1 Setting and Data processing
The Target Variable Construction
The focus of this study is the future change in monthly SST in the Nino3 region, which could be an important indicator of ENSO events. The target variable is manually crafted from the dataset using API and is visualized in Fig. 42, where the oscillation pattern can be seen.
1 # nino3 sst
2 y = Observation(
3 measurement=’Skin Temperature’,
4 level=’Surface’,
5 region={’latitude_min’: -5, ’latitude_max’: 5, ’longitude_min’: 210, ’longitude_max’: 270}
6 )
7
8 # next month shift
9 y = y.shift(1) - y
Listing 1: Code to define the target variable in ENSO case study
Figure 42: The visualization of the target variable defined in the ENSO case study. The oscillation pattern can be seen.
Prepare Tools for LLMs
The dataset is in the NetCDF format (network Common Data Format), so it is not convenient to be directly fed to LLMs. We prepare a function Observation(measurement, level, region) to help define factors. For example, in Lst. 1, we define the target variable with the help of this function. The measurement specifies the climate variables like precipitation rate or temperature; the level specifies the vertical location, like surface or specific pressure level; the region is a rectangle about the area concerned. This function would output a single time series about the measurement on the specified level averaged within the specified region. Note that the target variable is the only factor defined by humans, all other factors are defined by code from LLMs.
1 def get_air_temperature_at_cloud_layer_top_nino3():
2 ’’
3 Air Temperature at Cloud Layer Top (Nino3 Region)
4 ’’
5 # Declare observation
6 observation_name = Observation(
7 measurement="Air Temperature",
8 level="Isentropic Levels",
9 region={’latitude_min’: -5, ’latitude_max’: 5, ’longitude_min’: 210, ’longitude_max’: 270},
10 detailed_level=300
11 )
12
13 # Return the observation values directly
14 return observation_name
15
16 factor_dict[’Air Temperature at Cloud Layer Top (Nino3 Region)’] = get_air_temperature_at_cloud_layer_top_nino3()
Listing 2: LLM can be prompted to propose factors using complex tools
Feedback Construction
The dataset is in a special format; we can only draw a tabular subset of data from it in this setting. Therefore, it is difficult to make groups on samples like what we did in the previous cases. Instead, we construct feedback with some summary information. As shown in Fig. 43, two types of information are included in the feedback: intermediate causal graph and OLS regression result of factors in the current estimation of a Markov blanket on the target variable. These two types of information can be drawn naturally from the COAT’s intermediate results and are not complex, so LLMs can always access them and easily understand them. Intuitively, these factors would help LLMs better understand existing factors and propose further possible factors based on the provided dataset and its knowledge.
Figure 43: The prompt with feedback based on previous loop in the ENSO case study
K.2 Causal Graph and Discussion
Result Interpretation
In this case study, we are treating non-stationary climate data, while assuming the causal structure is invariant. Under this assumption, the non-stationarity is actually helpful for causal structure learning [51, 114]. To this end, we utilize the CD-NOD algorithm [51] to fully utilize the changing causal modules for better identifiability. CD-NOD would first identify the non-stationary nodes, whose conditional distribution given the causal parents are changing with time, and then use them for better causal structure recovery. Four factors are identified to be non-stationary, as shown in Fig. 44.
We have adjusted the nodes’ names, shapes, and colors for better visualization. There are 14 nodes in total, with 13 factors identified by COAT. Each factor is a time series about a certain climate measurement above a specific level averaged over a specific region. For simplicity, we only considered instantaneous causal relations among those time series.
There are three regions identified to be relevant to the ENSO phenomenon:
  • Equatorial Pacific Region (Orange Nodes). This region (5N-5S, 120W-280W) is one of the most active places about ENSO. It becomes significantly warm during the El Niño phase and becomes significantly cold during the La Niña phase [115].
  • Nino3 Region (Blue Nodes). This region (5N-5S, 150W-90W) is one of the most classical regions to monitor the El Niño events by scientists. It is also used by humans to construct the target variables [110].
  • South American Coastal Region (Green Nodes). This region (0N-20S, 80W-60W) includes the Peruvian coastal up-welling system, and is noticed to be relevant to the ENSO cycle [116].
Some insights are delivered by paths in the output causal graph:
[…]
  • Momentum Flux, V-compoent. This factor is about the vertical movement of air. It is crucial in driving atmospheric convection [41], and it is related to the Walker Circulation, which is an important component in the ENSO system[42]. Also, it could influence the change in sea level pressure and indirectly influence the SST change.
  • Cloud Cover. The factor is the fraction of the sky covered by clouds in the NINO3 region. It could influence the SST Change through solar radiation as well as water circulation. It is confirmed to have a significant correlation [43] with ENSO events and plays an important role in the atmospheric circulation and hydrological cycle [44].
  • Soil Temperature. This might be a novel hypothesis proposed by COAT, since we found no sufficient research to confirm this point to the best of our knowledge. This causal graph has also suggested two possible indirect mechanisms: (1) through Volumetric Soil Moisture and Cloud Cover; and (2) through Convective Precipitation Rate and Sea Level Pressure. Therefore, this finding encourages more serious investigations of these hypotheses.
Figure 44: The final causal graph found by COAT in the ENSO case study
Discussion on Additional assumption
The constructed feedback contains the causal structure among existing factors and their OLS regression analysis on the target variable. By doing so, some additional assumptions are implicitly made. For example, it assumes a linear relation between factors and the target variables. More detailed theoretical analysis is out of the scope of this case study, and we left it for future work.
Discussion on Hallucination
Although we have clarified how to use tools in prompts, LLM can sometimes propose unsupported data requests. For example, only certain pressure levels are supported in the dataset. To this end, we include a Python code in the prompt, as listed in Lst. 3. This function can check whether the LLM’s request is supported. We ask LLM to run this function in its code interpreter to make sure the proposed factors are valid.
1 import pandas
2
3 def check_choice(measurement, level, detailed_level=None):
4
5 meta = pd.read_csv(’Measurement_Level_meta.csv’)
6
7 measurement_set = set(meta.Measurement.values)
8 measurement_level_map = meta.groupby([’Measurement’]).agg({"Level": lambda x: set(x)})
9 detailed_level_range = {
10 ’Pressure Levels’: [1000, 975, 950, 925, 900, 850, 800, 750, 700, 650, 600, 550, 500, 450, 400, 350, 300, 250, 200, 150, 100, 70, 50, 30, 20, 10, 5, 1],
11 ’Multiple Subsurface Levels’:[0, 10, 40, 100],
12 ’Fixed Levels Above the Surface’: [12, 20, 30, 50, 80, 100, 150, 200, 250, 300, 500],
13 ’Isentropic Levels’: [300, 330, 350]
14 }
15
16 if measurement not in measurement_set:
17 return False, f"measurement error"
18 if level not in measurement_level_map.loc[measurement,’Level’]:
19 return False, f"level error, only allowed: {measurement_level_map.loc[measurement,’Level’]}"
20
21 if level in detailed_level_range.keys():
22 if detailed_level not in detailed_level_range[level]:
23 return False, f"need detailed level in {detailed_level_range[level]}"
24
25 return True, None
Listing 3: LLM is required to use this function to overcome hallucination
Appendix L Broader Impacts
This work focuses on fully leveraging the rich knowledge learned by LLMs during pre-training to facilitate causal discovery from unstructured data, with the hope of empowering broader applications and social benefits. Besides, this paper does not raise any ethical concerns. This study does not involve any human subjects, practices to data set releases, potentially harmful insights, methodologies and applications, potential conflicts of interest and sponsorship, discrimination/bias/fairness concerns, privacy and security issues, legal compliance, and research integrity issues.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See the end of the introduction part.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Sec. 2.5 and Appendix B.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: see Appendix D.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Appendix E.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: See Sec. 3.1, Sec. 4.1 and Appendix E.
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [Yes]
  33. Justification: Standard Deviations are included.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: See Appendix E.9.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: Authors follow the NeurIPS Code of Ethics.
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: See the Appendix L.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: No sensitive or risky content is included in this paper.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: See Appendix H.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [Yes]
  63. Justification: See Appendix H.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: not involve crowdsourcing nor research with human subjects.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.